# Exact predicates, exact constructions and combinatorics

### v1

@@ 2.2.1. The Weiler model
To evaluate boolean expressions, we construct a combinatorial representation of the relations within a volumetric mesh, called the Weiler model (Weiler, 1985). In what follows, I shall use the notations of combinatorial maps (Lienhardt, 1988).
A 3-dimensional combinatorial map (3-map) is defined as a quadruple $({\mathcal{D}},\sigma_{1},\alpha_{2},\alpha_{3})$, where ${\mathcal{D}}$ is a set of $N$ discrete elements called darts, symbolized as black arrows in Figure 9, that can be identified with the integers $[1..N]$. The three functions $\sigma_{1},\alpha_{2},\alpha_{3}$, acting on the set of darts ${\mathcal{D}}$, are defined as follows:
@@
  • $\alpha_{3}$ is an involution that maps each dart to the opposide dart within the adjacent surface sheet (triple red lines in Figure 9).
As can be seen, each triangle is composed of three darts. In practice, only triangles and $\alpha_{2},\alpha_{3}$ are stored explicitly. A triangle with index $t$ corresponds to three darts $3t$, $13t+$ 1, $23t+$ 2, and the involution $\sigma_{1}$ is given by $\sigma_{1}=d-(d\mod 3)+(d+1)\mod 3$.
@@
In addition, we define a notion of patch. The patch incident to a dart $d$ is defined by $<\sigma_{1},\bar{\alpha_{2}}>(d)$, where $\bar{\alpha_{2}}>(d)$ is defined by:
@@
Note that the zone where two shells are in contact corresponds to two different charts, one for each side.
@@ 2.2.2. Constructing the Weiler model
As for the predicates used in the other stages of the pipeine, these two predicates are filtered using interval arithmetics. The vector ${\bf p}_{2}-{\bf p}_{1}$ of the radial edge and the normal vector ${\bf n}_{0}$ are computed at the beginning of radial sort, in both interval and exact arithmetic.
Note that in these predicates, the two points ${\bf p}_{1}$ and ${\bf p}_{2}$ along the radial edge are intersection points, hence they are represented with homogeneous exact coordinates. As a consequence, even with the interval filters, sorting all the radial edges takes a significant amount of time. However, one can gain significant performance by noticing that the order around a radial edge remains the same when travering polygonal lines incident to the same set of patches. Hence one can propagate the order along radial polylines, stopping at the vertices incident to more than two bundles, shown in red in the small figure.
@@
  • The links $\sigma_{0}$ (that connect each dart to its successor around the triangle) and $\alpha_{3}$ (that connect each part of triangles adjacent to the same chart) are trivial to obtain (they do not depend on the radial sort);
@@ 2.2.3. Classifying
Once the Weiler model is constructed, the next step consists in classifying the triangles (step 3 in the processing pipeline shown in Figure 2 on page 2).
@@
  • compute $I(d)$. The connected component may be an internal boundary (for instance if you compute the difference between two concenric balls). It could be also included inside another object. It is not possible to deduce $I(d)$ from the sole combinatorial information, it requires some geometric tests: first, $I(d)$ is initialized to $\emptyset$. Then, a ray is launched from a point in $t(d)$, and the inside/outside status of the operators $B(t)$ arre flipped for each intersected triangle $t$ (again, in terms of bit manipulation, this simply means XOR-ing $I(d)$ with $B(t)$.
@@ 2.4. Nuts and bolts: two arithmetic kernels
I shall now give more details about the arithmetic kernels. The reader may skip this section in a first read. This section contains details that I considered worth sharing, about how to implement arbitrary precision numbers and about how to derive the formulas for the predicates acting on them, together with non-trivial details that have important consequence on performance. I tested two different version, one based on arithmetic expansions, and the other one based on multiprecision floating point arithmetics. Each kernel provides an exact_nt number type, that supports addition, subtraction and product, 2d and 3d vectors with cartesian and homogeneous coordinates represented by exact_nt, as well as the standard predicates orient_2d, orient_3d, and symbolically perturbed in_circle. Each kernel also provides lexicographic points comparison, used to create the global vertex table (see Section 2.1.3) and used by the symbolic perturbations. For both kernel, I detail how numbers are represented, and how the different predicates are implemented.
@@ 2.4.1. The expansion-based arithmetic kernel
Compression is an operation that takes an expansion and that returns a more compact expansion with the same value. It is described in Section 2.8 of (Shewchuk, 1997), and can be summarized as in the algorithm below888 I think there is a typo in the original article, line 14 should read $h_{top}\leftarrow q$ (small $q$ and not capital $Q$). I think that this error was overlooked before because compression is mostly needed when cascading operations, as done here since expansions are used in exact constructions, and it was probably not done before. .
@@
If you use $l_{i}=x_{i}^{2}+y_{i}^{2}$, you obtain exactly the same result as $\tt in\_circle$. Now, if you use instead $l_{i}=\mbox{round\_to\_neareset}(x_{i}^{2}+y_{i}^{2})$, then what you obtain is the same result as $\tt in\_circle\_weighted$, with $\psi_{i}=\mbox{round\_to\_neareset}(x_{i}^{2}+y_{i}^{2})-(x_{i}^{2}+y_{i}^{2})$. While it is not exactly the same result as $\tt in\_circle$, what you obtain in the end is still a well-defined object, a Regular Triangulation (dual of power diagram), very similar to the Delaunay triangulation, except for a few flipped edges, and more importantly, provided that always the same $l$ is used for the same point, and with adapted symbolic perturnation, the triangulation remains uniquely defined. There is however an important difference: if in_circle_l$({\bf p}_{1},{\bf p}_{2},{\bf p}_{3},{\bf p}_{4},l_{1},l_{2},l_{3},l_{4})$ is negative, it does not necessarily imply that $({\bf p}_{1},{\bf p}_{2},{\bf p}_{3},{\bf p}_{4})$ forms a convex quadrilateral. In the constrained Delaunay triangulation, this condition needs to be explicitly tested before flipping an edge.
@@
However, one still reaches the limit when attempting to create the Weiler model: the involved predicates compute cross-products between vectors joining constructed points, and then dot product between those. Remember that the points are themselves the result of intersections. I instrumented the code to output an histogram of the lengths of the expansions, and they can be as long as 65000. This is not that surprising, knowing that each operation can potentially double the length of the expansions, this corresponds to 16 nested levels. Thanks to compression, this seldom occurs (no more than a few times in multi-million element meshes). Besides the time and space requirement for these very long expansions, a more important problem is that they can yield overflows and underflows when computing products with them. For instance, using the expansion-based kernel, it can sucessfully evaluate the CSG trees in example0001.csg to example0020.csg in the OpenSCAD category of the ThingiCSG testsuite (see 3.2 below), but it fails with all CSG trees between example0021.csg to example0024.csg.
@@ 2.4.2. The multiprecision floating-point arithmetic kernel
For this reason, and because the algorithm is used in production by the Tessael company, I implemented and tested an alternative kernel, that does not have the limitations of the expansion-based kernel mentioned above.
@@
With this convention, the representation of a point is unique, and one can simply use the lexicographic order on $x,y,z,w$, forgetting their geometric nature. Clearly, it will give a different order as compared to the lexicographic order on $x/w,y/w,z/w$ used before, but it is not a problem since the only thing we need to have for the symbolic perturbation and for the global vertex table to work is a total order. We also noticed that we are not obliged to pre-shift the numbers so that $w$’s exponent is zero, instead of that we pass the shifts to the comparison function, and shift by their difference only. It makes both the spatial indexing and predicates significantly more efficient (else they keep shifting the same numbers left and right).
@@ 3.1. Thingi10K
I shall now report some timings and statistics, first with Thingi10K (Zhou and Jacobson, 2016), a database with 10000 meshes. In the database, 4523 of them have intersections. The complete subset of models with intersections is processed in 1h 45min. One of them (996816, shown in Figure 14) is particular and takes 1271s (20 min) to be processed. It has up to several thousands intersections in the same triangle. This is because it has in a tiny zone a “3D grid of triangles”, that is, exactly the best method to create $O(N^{3})$ intersections with $N$ triangles. It is a good crash test for all the components of a mesh intersection algorithm, in particular the constrained Delaunay triangulation. Timings for the next 10 meshes with the largest number of intersections are reported in Table 1. The first column gives the timings for the state of the art (Cherchi et al., 2020). The second column corresponds to the kernel based on arithmetic expansions. For one of the models (101633, displayed in Figure 15), this kernel could not compute a correct result, because underflows were encountered. It is explained by the shape of the triangles that have intersections, that are very skinny and intersecting near the pole. They generate arithmetic operations that combine very large and very small numbers. This also explains the long timing obtains with previous work on this model. The third column reports the timings obtained with the multi-precision floating point kernel. As can be seen, for the largest models, timings are faster than with (Cherchi et al., 2020), and for some of them they are slower. Faster timings are explained by our exact constructions: in a certain sense, indirect predicates need to redo the same computations several times, whereas exact constructions act as a “cache”. Slower timings are explained by the constrained Delaunay triangulations that I compute, that involve the rather costly in_circle predicate. As far as 101633 is concerned, carefully designing the arithmetic kernel and the associated predicates as done in Section 2.4.2 has a significant impact on the performance. The fourth column reports the timings obtained with the multi-precision floating point kernel and constrained triangulations (not Delaunay). It lets us see much it costs to ensure the Delaunay property. With this kernel, timings are almost always faster as compared to previous work. Note that we loose the uniqueness of the triangulation, and therefore one would need a pocket identification algorithm as in (Cherchi et al., 2020) to make a fair comparison.

### v2

@@ 2.2.1. The Weiler model
To evaluate boolean expressions, we construct a combinatorial representation of the relations within a volumetric mesh, called the Weiler model (Weiler, 1985). In what follows, I shall use the notations of combinatorial maps (Lienhardt, 1988). This notion was introduced for writing the algebraic specification of geometric algorithm operating on meshes. In our case it is useful as a way of compacly writing both the algorithmic description of the method and its computer implementation. In a combinatorial map, a mesh is represented as a set of objects (called darts) and connections between them. Combinatorial maps come also with higher level operations to manipulate and to navigate in portions of the mesh. Here we use a 3-dimensional combinatorial map (3-map) to represent the Weiler model (or 3d mesh arrangement). Intuitively, in a 3-map, the combinatorial elements (darts) are very similar to the classical halfedges used in surface meshes, with additional volumetric links that connect the boundaries of adjacent volumes (Fig. 9). Zones where two volumes touch each other are represented as “twin sheets” of darts with opposite orientations connected with volumetric links.
Formally, a 3-dimensional combinatorial map (3-map) is defined as a quadruple $({\mathcal{D}},\sigma_{1},\alpha_{2},\alpha_{3})$, where ${\mathcal{D}}$ is a set of $N$ discrete elements called darts, symbolized as black arrows in Figure 9, that can be identified with the integers $[1..N]$. The three functions $\sigma_{1},\alpha_{2},\alpha_{3}$, acting on the set of darts ${\mathcal{D}}$, are defined as follows:
@@
  • $\alpha_{3}$ is an involution that maps each dart to the opposite dart within the “twin” surface sheet (triple red lines in Figure 9).
As can be seen, each triangle is composed of three darts. In practice, only triangles and $\alpha_{2},\alpha_{3}$ are stored explicitly. A triangle with index $t$ corresponds to three darts $3t$, $13t+$ 1, $23t+$ 2, and the permutation $\sigma_{1}$ is given by $\sigma_{1}=d-(d\mod 3)+(d+1)\mod 3$.
@@
In addition, we define a notion of patch. The patch incident to a dart $d$ is defined by $<\sigma_{1},\bar{\alpha_{2}}>(d)$, where $\bar{\alpha_{2}}(d)$ is defined by:
@@
Note that the zone where two shells are in contact corresponds to two different patches, one for each side.
@@ 2.2.2. Constructing the Weiler model
As for the predicates used in the other stages of the pipeline, these two predicates are filtered using interval arithmetics. The vector ${\bf p}_{2}-{\bf p}_{1}$ of the radial edge and the normal vector ${\bf n}_{0}$ are computed at the beginning of radial sort, in both interval and exact arithmetic.
Note that in these predicates, the two points ${\bf p}_{1}$ and ${\bf p}_{2}$ along the radial edge are intersection points, hence they are represented with homogeneous exact coordinates. As a consequence, even with the interval filters, sorting all the radial edges takes a significant amount of time. However, one can gain significant performance by noticing that the order around a radial edge remains the same when traversing polygonal lines incident to the same set of patches. Hence one can propagate the order along radial polylines, stopping at the vertices incident to more than two bundles, shown in red in the small figure.
@@
  • The links $\sigma_{0}$ (that connect each dart to its successor around the triangle) and $\alpha_{3}$ (that connect each part of triangles adjacent to the same patch) are trivial to obtain (they do not depend on the radial sort);
@@ 2.2.3. Classifying
Once the Weiler model is constructed, the next step consists in classifying the triangles (step 3 in the processing pipeline shown in Figure 2 on page 2). The method is very similar to what is done in (Zhou et al., 2016) (with the difference that the interface between two volumetric regions is represented by two sheets instead of a single sheet, making classification a little bit easier).
@@
  • compute $I(d)$. The connected component may be an internal boundary (for instance if you compute the difference between two concenric balls). It could be also included inside another object. It is not possible to deduce $I(d)$ from the sole combinatorial information, it requires some geometric tests: first, $I(d)$ is initialized to $\emptyset$. Then, a ray is launched from a point in $t(d)$, and the inside/outside status of the operators $B(t)$ arre flipped for each intersected triangle $t$ (again, in terms of bit manipulation, this simply means XOR-ing $I(d)$ with $B(t)$. In terms of predicates, ray-triangle intersection uses orient_3d, computed using the coordinates of the input points and the exact coordinates of the intersections, (more on this in (§2.4 p.2.4)), therefore the “ray leakage” phenomenon observed when using approximate coordinates cannot occur. However, the configuration where a ray traverses exactly an edge of the triangle requires special handling. I tested two approaches. The first one uses symbolic perturbation (Edelsbrunner and Mücke, 1994). The second one (less elegant but much simpler) keeps launching rays with random directions until no degeneracy is encountered. Similar performance was obtained (remember that only one ray launching per connected component of the volumetric mesh is required). I recommend the second approach (though less elegant, it is far simpler to implement).
@@ 2.4. Nuts and bolts: two arithmetic kernels
I shall now give more details about the arithmetic kernels. The reader may skip this section in a first read. This section contains details that I considered worth sharing, about how to implement arbitrary precision numbers and about how to derive the formulas for the predicates acting on them, together with non-trivial details that have important consequence on performance. I tested two different versions, one based on arithmetic expansions, and the other one based on multiprecision floating point arithmetics. Each kernel provides an exact_nt number type, that supports addition, subtraction and product, 2d and 3d vectors with cartesian and homogeneous coordinates represented by exact_nt, as well as the standard predicates orient_2d, orient_3d, and symbolically perturbed in_circle. Each kernel also provides lexicographic points comparison, used to create the global vertex table (see Section 2.1.3) and used by the symbolic perturbations. For both kernels, I detail how numbers are represented, and how the different predicates are implemented.
@@ 2.4.1. The expansion-based arithmetic kernel
Compression is an operation that takes an expansion and that returns a more compact expansion with the same value. It is described in Section 2.8 of (Shewchuk, 1997), and can be summarized as in the algorithm below999 I think there is a typo in the original article, line 14 should read $h_{top}\leftarrow q$ (small $q$ and not capital $Q$). I think that this error was overlooked before because compression is mostly needed when cascading operations, as done here since expansions are used in exact constructions, and it was probably not done before. .
@@
If you use $l_{i}=x_{i}^{2}+y_{i}^{2}$ (computed exactly), then you obtain exactly the same result as $\tt in\_circle$. Now, consider that you replace the (exact) value of $l_{i}$ with the nearest floating-point number, that is, $l_{i}=\mbox{round\_to\_neareset}(x_{i}^{2}+y_{i}^{2})$, then what you obtain is the same result as $\tt in\_circle\_weighted$, with $\psi_{i}=\mbox{round\_to\_neareset}(x_{i}^{2}+y_{i}^{2})-(x_{i}^{2}+y_{i}^{2})$. While it is not exactly the same result as $\tt in\_circle$, what you obtain in the end is still a well-defined object, a Regular Triangulation (dual of power diagram), very similar to the Delaunay triangulation, except for a few flipped edges, and more importantly, provided that always the same $l$ is used for the same point, and with adapted symbolic perturnation, the triangulation remains uniquely defined. There is however an important difference: if in_circle_l$({\bf p}_{1},{\bf p}_{2},{\bf p}_{3},{\bf p}_{4},l_{1},l_{2},l_{3},l_{4})$ is negative, it does not necessarily imply that $({\bf p}_{1},{\bf p}_{2},{\bf p}_{3},{\bf p}_{4})$ forms a convex quadrilateral. In the constrained Delaunay triangulation, this condition needs to be explicitly tested before flipping an edge.
@@
However, one still reaches the limit when attempting to create the Weiler model: the involved predicates compute cross-products between vectors joining constructed points, and then dot product between those. Remember that the points are themselves the result of intersections. I instrumented the code to output a histogram of the lengths of the expansions, and they can be as long as 65000. This is not that surprising, knowing that each operation can potentially double the length of the expansions, this corresponds to 16 nested levels. Thanks to compression, this seldom occurs (no more than a few times in multi-million element meshes). Besides the time and space requirement for these very long expansions, a more important problem is that they can yield overflows and underflows when computing products with them. For instance, using the expansion-based kernel, it can sucessfully evaluate the CSG trees in example0001.csg to example0020.csg in the OpenSCAD category of the ThingiCSG testsuite (see 3.3 below), but it fails with all CSG trees between example0021.csg to example0024.csg.
@@ 2.4.2. The multiprecision floating-point arithmetic kernel
For this reason, and because the algorithm is used in production by Yoyodine Corp101010to be replaced with the real company name in the final non-anonymous version, I implemented and tested an alternative kernel, that does not have the limitations of the expansion-based kernel mentioned above.
@@
With this convention, the representation of a point is unique, and one can simply use the lexicographic order on $x,y,z,w$, forgetting their geometric nature. Clearly, it will give a different order as compared to the lexicographic order on $x/w,y/w,z/w$ used before, but it is not a problem since the only thing we need to have for the symbolic perturbation and for the global vertex table to work is a total order. I also noticed that we are not obliged to pre-shift the numbers so that $w$’s exponent is zero, instead of that we pass the shifts to the comparison function, and shift by their difference only. It makes both the spatial indexing and predicates significantly more efficient (else they keep shifting the same numbers left and right).
@@ 3.1. Thingi10K
I shall now report some timings and statistics, first with Thingi10K (Zhou and Jacobson, 2016), a database with 10000 meshes. In the database, 4523 of them have intersections. The complete subset of models with intersections is processed in 1h 45min. One of them (996816, shown in Figure 14) is particular and takes 1271s (20 min) to be processed. It has up to several thousands intersections in the same triangle. This is because it has in a tiny zone a “3D grid of triangles”, that is, exactly the best method to create $O(N^{3})$ intersections with $N$ triangles. It is a good crash test for all the components of a mesh intersection algorithm, in particular the constrained Delaunay triangulation. Timings for the next 10 meshes with the largest number of intersections are reported in Table 1. The first column gives the timings for the state of the art (Cherchi et al., 2020). The second column corresponds to the kernel based on arithmetic expansions. For one of the models (101633, displayed in Figure 15), this kernel could not compute a correct result, because underflows were encountered. It is explained by the shape of the triangles that have intersections, that are very skinny and intersecting near the pole. They generate arithmetic operations that combine very large and very small numbers. This also explains the long timing obtains with previous work on this model. The third column reports the timings obtained with the multi-precision floating point kernel. As can be seen, for the largest models, timings are faster than with (Cherchi et al., 2020), and for some of them they are slower. Faster timings are explained by our exact constructions: in a certain sense, indirect predicates need to redo the same computations several times, whereas exact constructions act as a “cache”. Slower timings are explained by the constrained Delaunay triangulations that I compute, that involve the rather costly in_circle predicate. As far as 101633 is concerned, carefully designing the arithmetic kernel and the associated predicates as done in Section 2.4.2 has a significant impact on the performance. The fourth column reports the timings obtained with the multi-precision floating point kernel and constrained triangulations (not Delaunay). It lets us see much it costs to ensure the Delaunay property. With this kernel, timings are almost always faster as compared to previous work. Note that we loose the uniqueness of the triangulation, and therefore one would need a pocket identification algorithm as in (Cherchi et al., 2020) to make a fair comparison. To further compare both algorithms, that is, (Cherchi et al., 2020) and our method with constrained Delaunay triangulation, I conducted additional experiments, described in the next subsection.
Figure 16. Boolean operations with large triangle meshes and with highly degenerate configurations. The bottom one (E) is shown without and with simplification of co-planar facets. (F) shows the correct output of (D) on the left part, and the result obtained with Cherchi et.al’s method on the right part, with spurious facets.
[…]
Boolean operations with large triangle meshes and with highly degenerate configurations.
ID ours Zhou et.al Cherchi et.al
dragon-bunny(A) 2.7 3.9 1.3
buddhaUlion(B) 11.5 16.1 5.8 (sometimes crashes)
cylUcyl(C) 3.7 25.8 28.2
20rods-20rods(D) 0.9 3.3 2.1 (incorrect result)
Table 2. Comparison with (Zhou et al., 2016) and (Cherchi et al., 2022), timings in seconds. Meshes are shown in Figure 16.
ID AABB Box $\cap$ Box $\Delta\cap\Delta$ CDT Radial Weiler model Classification Misc. Total Simplify
construction sort combinatorics coplanar
dragon-bunny (A) 106 172 1420 162 2 302 241 295 2700 10
buddhaUlion (B) 512 811 6323 471 3 1969 10 1401 11500 24
20rods-20rods (C) ¡1 1 14 717 20 40 15 54 864 72
cylUcyl (D) 1 6 183 3429 ¡1 20 1 20 3660 57
rot_cube_20 (E) ¡1 ¡1 ¡1 646 23 36 1 3 712 15
Table 3. Timing breakdown (in milliseconds). Meshes are shown in Figure 16.
3.2. Comparison with Zhou et.al 2016 and Cherchi et.al 2022
In this section, I compare the new boolean operation algorithm with (Zhou et al., 2016) and (Cherchi et al., 2022) that employ a very similar algorithm. Typical examples are reported in Table 2 (the corresponding meshes are shown in Figure 16). Examples A and B (scanned meshes with many small triangles, in generic position) are taken from (Douze et al., 2015). Example C (union between two “camembert cheese” with many co-planar small triangles) is inspired by (Cherchi et al., 2022) (Fig. 3). Examples D and E are inspired by (Trettner et al., 2022) (Figs 24 and 25). For meshes with many tiny triangles in generic position, as in the dragon-bunny (A) and buddhaUlion (B) test cases, the new method is faster than (Zhou et al., 2016) and slower than (Cherchi et al., 2022). For meshes with many co-planar intersections, such as in the cylUcyl (C) and 20rods-20rods (D) test cases, the new method is up to $6x$ faster than previous work. Cherchi et.al’s method crashes unpredictably on the buddhaUlion test case, and produces for 20rods-20rods an incorrect result, with spurious hanging facets (shown in red in the right part of Fig. 16-F, the left part corresponds to the expected result).
In Table 3, I report the timing breakdown of the algorithm executed on the same datasets, where the different columns of the table correspond to:
  • AABB construction: construction of the Axis-aligned bounding box tree (§2.1.1 p.2.1.1);
  • Box $\cap$ Box: determinate candidate intersecting facets by traversing the AABB (§2.1.1 p.2.1.1);
  • $\Delta\cap\Delta$: determine triangle pairs intersections (§2.1.2 p.2.1.2);
  • CDT: Constrained Delaunay Triangulation (§2.1.3 p.2.1.3);
  • Radial sort: sort triangles around non-manifold edges (§2.2.1 p.2.2.1);
  • Weiler combinatorics: combinatorial part of Weiler model construction (§2.2.1 p.2.2.1);
  • Classification: find the boundary of the result from the co-refinement (§2.2.3 p.2.2.3);
  • Misc: this regroups detecting co-located vertices and flat facets in the input, re-ordering mesh elements for better locality and faster multithreading, and detecting intersections located exactly on input vertices;
  • Simplify coplanar: optional simplification of co-planar facets (§2.2.4 p.2.2.4).
As can be seen, the highly-optimized triangulation algorithm introduced here takes a minimal amount of time (more consequent for D that has many co-planar intersections, but still 6x faster than previous work). This is explained by both the carefully written predicates and the Constrained Delaunay Triangulation method that exploits the combinatorial information to minimize invoking the predicates and that minimizes dynamic memory allocation, making it especially efficient in a multithreading context. The algorithm makes a maximum use of the combinatorial information, hence radial sorting is done a limited number of times, as in (Zhou et al., 2016), and raytracing is only required once per connected component of the 3-Map to determine inclusion, in contrast with (Cherchi et al., 2022) that uses one ray-tracing query per surface patch. This comes at a significant price for creating the combinatorial structure, which takes up to 10% of the total computation time. Optionally it is possible to merge co-planar facets, as shown in Figure 16-E (see also Figure 11 in (§2.2.4 p.2.2.4) and Figure 18). It introduces a negligible overhead, thanks one again to the optimized CDT. Removing unnecessary vertices in flat zones is crucial when evaluating deep CSG trees, as in the next subsection. Without this post-processing, the number of triangles would quickly explode when chaining boolean operations. The different contributions and specific choices that I made gain important performance for boolean operations appearing in CAD-like CSG trees, that nearly systematically comprise highly degenerate configurations with many co-planar intersecting triangles. However, even in this type of configuration, the integer-based method EMBER is still spectacularly more efficient: according to the timings reported in (Trettner et al., 2022). On the 20rods-20rods (Fig. 24 in the EMBER article) it takes no more than 4.5 ms, and on a configuration similar to the rot_cube_20 (Fig. 25 in the EMBER article) it takes 5.9 ms.